OpenStudy (anonymous):

y-5=3(x-2) in slope intercept form is y=3x+11 am i right?

jimthompson5910 (jim_thompson5910):

y-5=3(x-2) y-5=3x-6 try it again

jimthompson5910 (jim_thompson5910):

you probably had 3x+6, which is incorrect